OpenStudy (anonymous):

How many times will the graph of y = x2 intersect the graph of y = –x2?

OpenStudy (anonymous):

\[ x^2=-x^2 \]Solve for \(x\).

OpenStudy (anonymous):

Basically, we want \[ (x,y_1) = (x,y_2) \]And since \(x=x\) is always true, we need \(y_1=y_2\), which in this case is:\[ x^2=-x^2 \]Do you know how to solve this?

OpenStudy (goformit100):

▬ ▬ Sir/Ma'am, A Warm Welcome To Open Study ▬ ▬